# Anisotropy of emergent large-scale dynamics in forced stratified shear flows

### v1

@@ 3 Results
This study considers forced stratified shear flows at $\mathrm{Pr}=1$, $\mathrm{Re}_{0}=50$, $\mathrm{Ri}_{0}=0.0125$, and $t_{\textrm{r}}=100$ in domains of $L_{z}=48$. These parameters, as will be shown in more detail below, are associated with the minimum value of the initial gradient Richardson number $\mathrm{Ri}_{\textrm{g,0}}(z=0)=\mathrm{Ri}_{0}R_{0}$, which is sufficiently small in principle to allow the development of primary KHI. Note that the initial gradient Richardson number $\mathrm{Ri}_{\textrm{g,0}}(z)$ is defined in terms of the initial profile,
$\mathrm{Ri}_{\textrm{g,0}}(z):=\mathrm{Ri}_{0}\frac{\frac{\partial b_{0}}{\partial z}}{\left(\frac{\partial u_{x,0}}{\partial z}\right)^{2}}.$ (11)
We investigate and quantify the emergent dynamics of the flow while varying the horizontal extents of the domain, $L_{x}$ and $L_{y}$. Normally, we consider domains of square horizontal cross-section with $L_{\mathrm{h}}\equiv L_{x}=L_{y}$ ranging from $16$ to $512$. Table 1 summarises all our considered domains.
@@ 3.1 Typical evolution of a forced stratified shear flow
Figure 2: Temporal evolution of forced, stratified shear flows. (a – e) The flow is prone to a primary KHI, leading eventually to ‘overturning billows’ and streamwise mergers. A continuous forcing sustains the induced turbulence for arbitrarily long times. During this evolution of the flow, (f – h) the interface broadens before reaching a statistically stationary depth. Note that for this relatively small $\mathrm{Re}_{0}$, as shown in (g), diffusion of the shear layer and density interface dominates the development of the primary instability until the shear layer and density interface have approximately doubled in depth. In this figure, $L_{\mathrm{h}}=128$ while panels (a – e) visualise $b\left(x,y=0,z,t\right)$ with the colour bar matching figure 6 (l, p).
We first consider the evolution of a typical flow in a domain with horizontal extent of $L_{\mathrm{h}}=128$. Figures 2(a – e) show snapshots of the instantaneous buoyancy field $b\left(x,y=0,z,t\right)$ in vertical slices normal to the spanwise direction. At early times, the flow is prone to a primary Kelvin-Helmholtz instability which, driven by the vertical shear, develops into a train of KH billows (panel (c)) that merge subsequently (panel (d)). These primary billows break down, and the buoyancy interface broadens (panel (f)). Note that we use $\langle\cdot\rangle_{\Phi}$ to denote averages over $\Phi$.
@@
Across the evolution of the flow, we quantify the associated broadening of the shear layer half-depth and buoyancy interface half-thickness via
[…]
respectively, using the approach proposed by (Smyth & Moum, 2000). Together with the initial profiles defined in equations (6) and (7), these functions yield $d\left(t=0\right)=1$ and $\delta\left(t=0\right)=1/R_{0}$ (as these non-dimensional lengths $\left\{d,\delta\right\}$ are measured in units of $d_{0}$, see again equation (1)). Similarly, the time-dependent ratio of interface (half) thicknesses $R\left(t\right):=d/\delta$ with $R\left(t=0\right)=R_{0}$. Figures 2 (g, h) highlight that both $d$ and $\delta$ converge after an initial transient to statistically stationary values $\left\{d,\delta\right\}\gg 1$. In other words, the interfaces have deepened significantly. Their ratio $R\approx 1$ at late times, consistently with the fact that the flow remains ‘weakly’ stratified. Here, ‘weakly stratified’ is meant in the specific sense that the turbulent diffusivity of buoyancy closely follows the turbulent diffusivity of the momentum. Equivalently, the turbulent Prandtl number is close to one, and so the buoyancy field is at least in some sense slaved to the velocity field and behaves like a passive scalar. We note in passing that, as highlighted in panel (g), the onset of the primary KHI (evidenced by the significant change in the rate of the growth of $d$ and $\delta$ around $t\simeq 150$) occurs only after a period of significant diffusive deepening of both the shear layer and the buoyancy interface. As the shear layer depth $d$ approximately doubles, the effective Reynolds number (see below) approximately doubles as well, thus eventually triggering the onset of the primary instability with a significantly larger wavelength than that expected when defined in terms of the initial shear layer half-depth $d_{0}$. Investigation of this interesting early-time interaction between diffusion and instability onset is beyond the scope of this paper but undoubtedly worthy of further, more detailed consideration.
Moreover, the above definition of $d\left(t\right)$ enables the computation of time-dependent values of the bulk Reynolds and Richardson number via
@@
Table 1 lists the temporal averages and standard deviations of $d$, $\delta$, $R$, $\mathrm{Re}$, and $\mathrm{Ri}$ during the late statistically stationary dynamics of all our simulations. We discuss their trends with respect to the horizontal domain size $L_{\mathrm{h}}$ in section 3.3.
3.2 Dependence of horizontally-averaged fields on domain size
@@
Figure 3: Evolution of the horizontally-averaged buoyancy field. (a) The deepening of the density interface is affected by the horizontal extent of the domain $L_{\mathrm{h}}\equiv L_{x}=L_{y}$ and converges at large values of $L_{\mathrm{h}}$ only. For $L_{\mathrm{h}}=128$: (b) the late temporal average of $\mathrm{Ri}_{\textrm{g}}$ shows a significant increase in its minimum value to $0.74$, with an associated increase throughout the turbulent ‘mixing zone’ (the canonical marginal value of $0.25$ is marked with a vertical dot-dashed line), while the resulting associated mixing in the mixing zone is underlined by high amplitudes in (c) the stabilising vertical buoyancy advection (i.e. the buoyancy flux) $B$ and (d) the dissipation rates of kinetic energy and scaled buoyancy variance, $\varepsilon_{u}$ and $\chi$.
@@
Turning attention to domain size effects, in figure 3 (a) we plot the mean vertical buoyancy profiles $\langle b\rangle_{A,t}$ associated with the late statistically stationary dynamics. We find that these profiles depend strongly on the horizontal extent $L_{\mathrm{h}}$ of the domain. Although $L_{\mathrm{h}}=16\gg 1$ even for our smallest domain, the interface deepens for increasing $L_{\mathrm{h}}$ and a convergence seems to be reached for $L_{\mathrm{h}}\gtrsim 64$ only. This suggests that the vertical mixing of buoyancy strongly depends on the horizontal extent of the domain.
@@
This observed deepening of the buoyancy interface is accompanied by a deepening of the shear layer, both affecting in turn the (time-dependent) gradient Richardson number
@@
$\mathrm{Ri}_{\textrm{g}}\left(z,t\right):=\frac{\langle N^{2}\rangle_{A}}{\langle S\rangle_{A}^{2}}$ (14)
@@
via the vertical shear $S=\partial u_{x}/\partial z$ and buoyancy frequency
$N:=\sqrt{\mathrm{Ri}_{0}\frac{\partial b}{\partial z}}.$ (15)
@@
We remark that $\mathrm{Ri}_{\textrm{g}}\left(z,t=0\right)=\mathrm{Ri}_{\textrm{g,0}}$ as defined in (11). As shown in figure 3(b), the minimum value of $\mathrm{Ri}_{\textrm{g}}$ has significantly increased (to $\langle\mathrm{Ri}_{\textrm{g}}\left(z=0\right)\rangle_{t}\approx 0.074$ for $L_{\mathrm{h}}=128$). Note that, due to the square in the denominator of the definition, the reduction of shear overpowers the simultaneous reduction of buoyancy stratification.
@@
From this panel (b), it is also apparent that $\mathrm{Ri}_{\textrm{g}}\leq 0.25$ within the region $-\langle d\rangle_{t}\leq z\leq\langle d\rangle_{t}$. Although the flow is undoubtedly turbulent – and so the Miles-Howard criterion (Miles, 1961; Howard, 1961) of (steady) linear inviscid stability theory is definitely not applicable –, such relatively small values of $\mathrm{Ri}_{\textrm{g}}$ are necessary for turbulence to survive and so we refer to this zone as the turbulent mixing zone.
@@
as well as the dissipation rates of kinetic energy and scaled buoyancy variance
@@
respectively. These (non-dimensional) quantities are related to the evolution equations of kinetic energy and buoyancy variance as shown in appendix A. Here, $\bm{S}$ represents the strain rate tensor and $\varepsilon_{b}$ the (unscaled) buoyancy variance dissipation rate. Note that the scaling via $\langle N^{2}\rangle_{A,t}$ results in identical physical units for the associated dimensional dissipation rates. We emphasise this point by the comparison
@@
where we have, for improved clarity, re-introduced tildes for non-dimensional quantities in the above three equations only. This property of identical physical units is particularly helpful when studying the exchange of kinetic and potential energy, as buoyancy variance is closely related to the concept of ‘available potential energy’ (Caulfield, 2021).
As shown in panels (c – d) of figure 3, all of these quantities exhibit pronounced peaks close to the midplane. However, while the vertical buoyancy advection $B$ introduces a macroscopic stirring that is generally reversible, due to the inherent coupling of $\bm{u}$ and $b$ it leads to irreversible dissipation via both $\varepsilon_{u}$ and $\varepsilon_{b}$ (Caulfield, 2021), see again appendix A. As $\langle B\rangle_{V,t}<0$ (with the domain volume $V:=A\times L_{z}$), this stirring comes with an overall stabilising effect on the configuration. Furthermore, although enhanced values of $\varepsilon_{u}$ and $\chi$ extend beyond the ‘mixing zone’, it is clear that the mixing zone contains the vast majority of the enhanced dissipation at these late times.
@@ 3.3 Convergence of vertical stirring, dissipation and mixing for extended domains
Figure 4: Impact of the horizontal extent of the domain on the mixing. The flow topology – comprising the (a) interface depths, (b) final emergent (bulk) Reynolds number and Richardson number, as well as averages across the midplane and across the entire mixing zone of (c) gradient Richardson number, (d) buoyancy flux, (e) kinetic energy dissipation and (f) buoyancy variance dissipation – only converges for horizontally extended domains, $L_{\mathrm{h}}\gtrsim L_{\mathrm{h,crit}}=96$. Vertical solid lines indicate the temporal standard deviation.
@@
After having introduced important quantities related to the mixing of the buoyancy field in the previous sections 3.1 and 3.2, in figure 4 we plot their variation as a function of $L_{\mathrm{h}}$. Underlining the conclusions from figure 3 (a), figure 4(a) shows that the (half-) thicknesses of the shear layer or the streamwise velocity and buoyancy interfaces converge for extended horizontal domains only. As the parameters $\left\{\mathrm{Re},\mathrm{Ri}\right\}\propto d$, they also only converge for extended domains, as shown in panel (b). Note both are significantly increased (effectively by a factor of eight) from their initial values. In panel (c) we also plot temporal averages of the gradient Richardson number $\mathrm{Ri}_{\textrm{g}}$ as defined in equation (14). Both our analyses for the midplane and the entire mixing zone demonstrate again a convergence for extended domains only. Unsurprisingly, we observe a similar behaviour for the mixing-related quantities $B$, $\varepsilon_{u}$, and $\chi$ as shown in panels (d – f).
@@
In summary, we find that the properties of the (predominantly vertical) stirring, dissipation and mixing actually depend strongly on the horizontal extent of the domain $L_{\mathrm{h}}$. The associated emergent vertical dynamics of the flow seems to become independent of $L_{\mathrm{h}}$ only once $L_{\mathrm{h}}\gtrsim L_{\mathrm{h,crit}}=96$ where $L_{\mathrm{h,crit}}$ is a critical extent of the domain.
@@
Figure 5: Statistics of mixing properties at midplane. Statistical distributions of (a) kinetic energy dissipation rate $\varepsilon_{u}$, (b) scaled buoyancy variance dissipation rate $\chi$, and (c) local flux coefficient $\Gamma_{\chi}$ are affected by insufficient horizontal extents of the domain. The grey dashed vertical line marks the canonical value $\Gamma_{\chi}=0.2$. Note the emergent scaling for $\Gamma_{\chi}$ for extreme mixing events, and the marked difference of the high tails of the PDFs of $\varepsilon_{u}$ and $\chi$.
@@
As shown in figure 5, the dependence of dissipation and mixing rates on $L_{\mathrm{h}}$ does not just affect mean values but even the general structure of their probability density functions (PDFs). Interestingly, we observe a pronounced scaling of the PDF of the local flux coefficient $\Gamma_{\chi}$ – defined as
$\Gamma_{\chi}:=\frac{\chi}{\varepsilon_{u}},$ (22)
@@
and representing the local ratio between dissipation of scaled buoyancy variance and kinetic energy – for values $\Gamma_{\chi}\gtrsim 0.2$, i.e., the canonical value proposed by Osborn (1980), as indicated in panel (c).
Indeed, as is apparent from figure 4 (e, f), the converged values of ‘bulk’ turbulent mixing coefficients
$\bar{\Gamma}_{\chi,\Phi}:=\frac{\langle\chi\rangle_{\Phi,t}}{\langle\varepsilon_{u}\rangle_{\Phi,t}},$ (23)
where the average is taken across either $z=0$ or $-\langle d\rangle_{t}\leq z\leq\langle d\rangle_{t}$, are substantially smaller than the canonical maximum value of $0.2$. We find that $\bar{\Gamma}_{\Phi}\simeq 0.079-0.081$ at the midplane and $\bar{\Gamma}_{\Phi}\simeq 0.13-0.14$ within the mixing zone, essentially because $\varepsilon_{u}$ is more localized at the midplane than $\chi$ is. This discrepancy relative to the proposal of Osborn (1980) is not entirely surprising, as it is clear that though both $\varepsilon_{u}$ and $\chi$ are elevated in the mixing zone, the assumption that they are linearly related (which is at the heart of Osborn’s influential arguments) is clearly not qualitatively or quantitatively correct.

### v2

@@ 3 Results
This study considers forced stratified shear flows at $\mathrm{Pr}=1$, $\mathrm{Re}_{0}=50$, $\mathrm{Ri}_{0}=0.0125$, and $t_{\textrm{r}}=100$ in domains of $L_{z}=48$. As will be shown in more detail in section 3.2, these parameters are associated with the minimum value of the initial gradient Richardson number $\mathrm{Ri}_{\textrm{g,0}}\left(z=0\right)=\mathrm{Ri}_{0}R_{0}$ which is sufficiently small to enable the development of primary KHI. We investigate and quantify the emergent dynamics of the flow while varying the horizontal extents of the domain, $L_{x}$ and $L_{y}$. Normally, we consider domains of square horizontal cross-section with $L_{\mathrm{h}}\equiv L_{x}=L_{y}$ ranging from $16$ to $512$. Table 1 summarises all our considered domains.
@@ 3.1 Typical evolution of a forced stratified shear flow
Figure 2: Temporal evolution of forced, stratified shear flows. (a – e) The flow is prone to a primary KHI, leading eventually to ‘overturning billows’ and streamwise mergers. A continuous forcing sustains the induced turbulence for arbitrarily long times. During this evolution of the flow, (f – h) the interface broadens before reaching a statistically stationary depth. Note that for this relatively small $\mathrm{Re}_{0}$, as shown in (g), molecular diffusion of the shear layer and density interface dominates the development of the primary instability until the depths of the shear layer and density interface have approximately doubled. In this figure, $L_{\mathrm{h}}=128$ while panels (a – e) visualise $b\left(x,y=0,z,t\right)$ with the colour bar matching figure 6 (l, p).
We first consider the evolution of a typical flow in a domain with horizontal extent of $L_{\mathrm{h}}=128$. Figures 2 (a – e) show snapshots of the instantaneous buoyancy field $b\left(x,y=0,z,t\right)$ in vertical slices normal to the spanwise direction. At early times, the flow is prone to a primary Kelvin-Helmholtz instability which, driven by the vertical shear, develops into a train of KH billows (panel (c)) that merge subsequently (panel (d)). These primary billows break down, and the buoyancy interface broadens (panel (f)). Note that we use $\langle f\rangle_{\Phi}$ to denote the average value of some flow variable $f$ computed over $\Phi$, where the averaging domain $\Phi$ may be a combination of spatial and temporal dimensions. Here, time-averages cover the flows’ last $\Delta t=3,000$, i.e. the time interval associated with the statistically stationary regime of primary interest.
@@
Across the evolution of the flow, we quantify the associated broadening of the shear layer and buoyancy interface via
$d\left(t\right):=\frac{1}{2}\int_{-L_{z}/2}^{+L_{z}/2}\left(1-\langle u_{x}\rangle_{A}^{2}\right)dz\qquad\textrm{and}\qquad\delta\left(t\right):=\frac{1}{2}\int_{-L_{z}/2}^{+L_{z}/2}\left(1-\langle b\rangle_{A}^{2}\right)dz,$ (11)
respectively, using the approach proposed by Smyth & Moum (2000). Together with the initial profiles defined in equations (6) and (7), these functions yield $d\left(t=0\right)=1$ and $\delta\left(t=0\right)=1/R_{0}$ (as these non-dimensional lengths $\left\{d,\delta\right\}$ are measured in units of $d_{0}$, see again equation (1)). Similarly, the time-dependent ratio of interface depths $R\left(t\right):=d/\delta$ with $R\left(t=0\right)=R_{0}$. Figures 2 (g, h) highlight that both $d$ and $\delta$ converge after an initial transient of length $\lesssim 10^{3}\tau_{\textrm{adv}}$ to statistically stationary values $\left\{d,\delta\right\}\gg 1$. In other words, the interfaces have deepened significantly and, remarkably, reached a finite value characterised by a balance between the ongoing forcing (attempting to relax the interfaces back to their original depths) and mixing (tending to deepen the interfaces). We establish the independence of this finite value from the extent of the domain in the subsequent sub-sections and hypothesise a physical reason for this convergence in section 4. Consistently with the fact that the flow remains ‘weakly’ stratified, the ratio $R\approx 1$ at late times. Here, ‘weakly stratified’ is meant in the specific sense that the turbulent diffusivity of buoyancy closely follows the turbulent diffusivity of the momentum. Equivalently, the turbulent Prandtl number is close to one, and so the buoyancy field is at least in some sense slaved to the velocity field and behaves like a passive scalar (Turner, 1973). Due to the particular form of the forcing, with shear continually being reinforced, it is also appropriate to consider this flow to be ‘shear-dominated’ within the nomenclature of Mater & Venayagamoorthy (2014).
We note in passing that, as highlighted in panel (g), the onset of the primary KHI (evidenced by the significant change in the rate of the growth of $d$ and $\delta$ around $t\simeq 150$) appears only after a period of significant diffusive deepening of both the shear layer and the buoyancy interface. As the shear layer half-depth $d$ approximately doubles during this initial period of time, the onset of the primary instability triggers a significantly larger wavelength. In other words, the characteristic scales of the delayed KHI are set relative to $d$ rather than $d_{0}$. An investigation of this interesting early-time interaction between diffusion and instability onset is beyond the scope of this paper but undoubtedly worthy of further, more detailed consideration.
Having reached the final, statistically stationary state, the above definition of $d\left(t\right)$ enables the computation of associated time-dependent values of the bulk Reynolds and Richardson number via
@@
Table 1 lists the temporal averages and standard deviations of $d$, $\delta$, $R$, $\mathrm{Re}$, and $\mathrm{Ri}$ during the late statistically stationary dynamics of all our simulations. We discuss their trends with respect to the horizontal domain size $L_{\mathrm{h}}$ in section $3.$
3.2 Sustained turbulent interactions between the two layers
@@
Figure 3: Sustained turbulent interactions. (a) Despite statistical stationarity, the deepening of the density interface can be affected by the horizontal extent of the domain $L_{\mathrm{h}}\equiv L_{x}=L_{y}$ before it converges eventually. This deepening tends to stabilise the emergent flow, resulting in a significantly increased minimum value of $0.074$ of (b) the average late $\mathrm{Ri}_{\textrm{g}}$. Nevertheless, we find $\mathrm{Ri}_{\textrm{g}}\leq 0.25$ (with this canonical value being marked with a vertical line) throughout the turbulent ‘mixing zone’. The resulting associated mixing in this zone is underlined by high amplitudes in (c) the stabilising vertical buoyancy advection (i.e. the buoyancy flux) $B$ and (d) the dissipation rates of kinetic energy and scaled buoyancy variance, $\varepsilon_{u}$ and $\chi$.
@@
Even though a continuous forcing allows the existence of a statistically stationary regime, the associated sustained turbulent interactions between the fluid regions above and below the shear layer may yet be impacted by the extent of the domain. In figure 3 (a) we plot the mean vertical buoyancy profiles $\langle b\rangle_{A,t}$ associated with the late-time dynamics of the forced stratified shear flow. We find that these profiles depend strongly on the horizontal extent $L_{\mathrm{h}}$ of the domain. Although $L_{\mathrm{h}}=16\gg 1$ even for our smallest domain, the interface deepens for increasing $L_{\mathrm{h}}$ and a convergence seems to be reached for $L_{\mathrm{h}}\gtrsim 64$ only. This suggests that the vertical stirring and resultant mixing of buoyancy strongly depends on the horizontal extent of the domain.
@@
This observed deepening of the buoyancy interface is accompanied by a deepening of the shear layer, both affecting in turn the gradient Richardson number
@@
$\mathrm{Ri}_{\textrm{g}}\left(z\right):=\frac{\langle N^{2}\rangle_{A,t}}{\langle S\rangle_{A,t}^{2}}\equiv\left(\frac{t_{\textrm{S}}}{t_{\textrm{N}}}\right)^{2}$ (13)
@@
via the buoyancy frequency (for which averaging is typically associated with $\partial b/\partial z$) and vertical shear,
$N:=\sqrt{\mathrm{Ri}_{0}\frac{\partial b}{\partial z}}\qquad\textrm{and}\qquad S:=\frac{\partial u_{x}}{\partial z},$ (14)
@@
respectively. We remark that it is also possible to associate the initial profiles from equations (6) and (7) with an initial gradient Richardson number $\mathrm{Ri}_{\textrm{g}}\left(z,t=0\right)=\mathrm{Ri}_{\textrm{g,0}}:=\mathrm{Ri}_{0}\left(\partial b_{0}/\partial z\right)/\left(\partial u_{x,0}/\partial z\right)^{2}$. As shown in figure 3 (b), the minimum value of $\mathrm{Ri}_{\textrm{g}}$ has significantly increased from $\mathrm{Ri}_{\textrm{g,0}}\left(z=0\right)=0.0125$ to $\mathrm{Ri}_{\textrm{g}}\left(z=0\right)\approx 0.074$ provided $L_{\mathrm{h}}=128$. Note that, due to the square in the denominator of the definition, the reduction of shear overpowers the simultaneous reduction of buoyancy stratification. As indicated on the right of equation (13), $\mathrm{Ri}_{\textrm{g}}$ can also be interpreted as a ratio of time scales associated with the average shear and stratification, an observation we will discuss in detail in section 4.
@@
From this panel (b), it is also apparent that $\mathrm{Ri}_{\textrm{g}}\leq 0.25$ within the region $\langle d\rangle_{t}\leq z\leq\langle d\rangle_{t}$. Although the flow is undoubtedly turbulent – and so the Miles-Howard criterion (Miles, 1961; Howard, 1961) of (steady) linear inviscid stability theory is definitely not applicable –, such relatively small values of $\mathrm{Ri}_{\textrm{g}}$ are necessary for turbulence to survive and so we refer to this zone as the turbulent mixing zone.
@@
as well as the dissipation rates of kinetic energy $\varepsilon_{u}$ and scaled buoyancy variance $\chi$
@@
respectively, which naturally emerge in the evolution equations for kinetic energy and scaled buoyancy variance (see for example the review of Caulfield (2021) for a derivation). Here, $\bm{S}$ represents the strain rate tensor and $\varepsilon_{b}$ the (unscaled) buoyancy variance dissipation rate. Note that the scaling via $\langle N^{2}\rangle_{A,t}$ results in identical physical units for the associated dimensional dissipation rates. We emphasise this point by the comparison
@@
where we have, for improved clarity, re-introduced tildes for non-dimensional quantities in the above three equations only. This property of identical physical units is particularly helpful when studying the exchange of kinetic and potential energy, as buoyancy variance is closely related to the concept of ‘available potential energy’ as originally introduced by Lorenz (1955) and further developed by Winters et al. (1995), see for example the review of Caulfield (2021) for further discussion.
As shown in panels (c – d) of figure 3, all of these quantities exhibit pronounced peaks close to the midplane. However, while the vertical buoyancy advection $B$ introduces a macroscopic stirring that is generally reversible, it leads to irreversible dissipation via both $\varepsilon_{u}$ and $\varepsilon_{b}$ due to the inherent coupling of $\bm{u}$ and $b$. As $\langle B\rangle_{V,t}<0$, this stirring comes with an overall stabilising effect on the configuration. Furthermore, although enhanced values of $\varepsilon_{u}$ and $\chi$ extend beyond the mixing zone, it is clear that this mixing zone contains the vast majority of the enhanced dissipation in this regime.
Hence, the emergent turbulent shear half-depth $\langle d\rangle_{t}$ – as defined in equation (11) – represents an appropriate measure of the mixing zone as a region of strong interactions between the two fluid layers. We discuss further implications of these vertical profiles (including for values of the average flux coefficient) in more detail in section 4.
@@ 3.3 Convergence of vertical stirring, dissipation and mixing for extended domains
Figure 4: Impact of the horizontal extent of the domain on the mixing. The flow topology – comprising the (a) interface (half-) depths, (b) final emergent (bulk) Reynolds number and Richardson number, as well as averages across the midplane and across the entire mixing zone of (c) gradient Richardson number, (d) buoyancy flux, (e) kinetic energy dissipation and (f) buoyancy variance dissipation – only converges for horizontally extended domains, $L_{\mathrm{h}}\gtrsim L_{\mathrm{h,crit}}=96$. Vertical solid lines indicate the temporal standard deviation. All panels share the same abscissa.
@@
After having introduced important quantities related to the vertical interaction of the velocity and buoyancy field in the previous sections 3.1 and 3.2, in figure 4 we plot their variation as a function of $L_{\mathrm{h}}$. Underlining the conclusions from figure 3 (a), figure 4 (a) shows that the depths of the shear layer or the streamwise velocity and buoyancy interfaces indeed converge for sufficiently extended horizontal domains. As the parameters $\left\{\mathrm{Re},\mathrm{Ri}\right\}\propto d$, they converge for extended domains, too, as shown in panel (b). Note that both are significantly increased from their initial values. Starting with panel (c), we plot spatio-temporal averages of various quantities where the spatial averages are performed either across the two-dimensional horizontal midplane or across the volume of the entire turbulent mixing zone. Unsurprisingly, given the convergence of the depth of the mixing zone, we observe similar convergence for the gradient Richardson number $\mathrm{Ri}_{\textrm{g}}$ as well as the vertical stirring- and mixing-related quantities $B$, $\varepsilon_{u}$, and $\chi$ as shown in panels (c – f).
@@
In summary, we find that the properties of the (predominantly vertical) stirring, dissipation and mixing actually depend strongly on the horizontal extent of the domain $L_{\mathrm{h}}$. However, the associated emergent vertical dynamics of the flow seems indeed to become independent of $L_{\mathrm{h}}$ once $L_{\mathrm{h}}\gtrsim L_{\mathrm{h,crit}}=96$ where $L_{\mathrm{h,crit}}$ is a critical extent of the domain.
@@
Figure 5: Statistics of mixing properties at midplane. Statistical distributions of the (a) kinetic energy dissipation rate $\varepsilon_{u}$, (b) scaled buoyancy variance dissipation rate $\chi$, and (c) local flux coefficient $\Gamma_{\chi}$ are affected by insufficient extents of the domain but converge eventually. The grey dashed vertical line marks the canonical value $\Gamma_{\chi}=0.2$. Note the emergent scaling for $\Gamma_{\chi}$ for extreme mixing events, and the marked difference of the high tails of the PDFs of $\varepsilon_{u}$ and $\chi$.
@@
Finally, as shown in figure 5, we remark that not only the mean values of of dissipation and mixing rates depend on $L_{\mathrm{h}}$, but also the general structure of their probability density functions (PDFs) depends on $L_{\mathrm{h}}$ as well. Interestingly, we observe a pronounced scaling of the PDF of the (point-wise) local flux coefficient $\Gamma_{\chi}$ – defined as
$\Gamma_{\chi}:=\frac{\chi}{\varepsilon_{u}}$ (21)
@@
and representing the local ratio between dissipation of scaled buoyancy variance and kinetic energy – for values $\Gamma_{\chi}\gtrsim 0.2$, i.e., the canonical value proposed by Osborn (1980), as indicated in panel (c). We return to a more conventional definition of a (in some sense) global flux coefficient – based on average rather than local quantities – in section 4.